# FPBoost: Fully Parametric Gradient Boosting for Survival Analysis

### v2

@@ 1 Introduction
In order to model the survival function, traditional methods often rely on simplifying assumptions, such as the risk proportion between different subjects being constant over time (Cox, 1972). These assumptions allow the construction of survival functions from a small set of parameters that can be estimated with statistical methods. While suited for contexts with limited data, these simplifying assumptions constrain generalization in real-world scenarios (Katzman et al., 2018). Machine learning techniques have allowed to advance these models by incorporating decision trees and neural networks, significantly enhancing their ability to identify and learn non-linear interactions within high-dimensional features. However, most of these techniques still operate under certain constraints, such as time discretization in neural-based approaches (Kvamme & Borgan, 2021) or accelerated failure time in tree-based gradient boosting (Collett, 2023).
@@
In this context, we introduce Fully Parametric Gradient Boosting (FPBoost), a novel architecture designed to model hazard functions through the composition of multiple fully parametric hazard functions. Hazard functions are related to survival functions as they measure the instantaneous risk of a subject experiencing the event of interest. FPBoost combines the strengths of tree-based ensemble learning with gradient boosting (Friedman, 2001; Collett, 2023), offering a flexible model with robust generalization capabilities and minimal assumptions. Modeling hazard as a weighted sum of multiple, fully-parametric functions, referred to as heads, allows FPBoost to be trained by maximizing the full survival likelihood (Wang et al., 2019). This, in turn, removes the need for simplified assumptions such as partial likelihood (Cox, 1972; Katzman et al., 2018) or discrete losses (Kvamme & Borgan, 2021). Additionally, the continuous nature of the learned survival functions ensures a fine-grained estimation of the event distribution, without requiring interpolation techniques (Archetti et al., 2024). We theoretically show how this framework makes FPBoost a universal approximator of hazard functions, allowing it to learn in principle any target hazard, provided enough heads. Lastly, gradient-boosted trees applied to tabular data – the most common data format in survival applications – has proven to be still competitive against neural network techniques (Grinsztajn et al., 2022), as highlighted by our empirical results.
@@
We evaluate FPBoost in the right-censored, single-event setting, which is the most common application of survival analysis. The performance of FPBoost is benchmarked against state-of-the-art survival models, including both tree-based and neural network-based models (Chen et al., 2024). Performance metrics include the concordance index (Uno et al., 2011) to measure the discrimination capabilities of the model and the integrated Brier score (Graf et al., 1999), tailored for calibration. Our experiments demonstrate that FPBoost outperforms alternative models in both discrimination and calibration in the majority of cases and matches their performance when it does not.
@@ 2.1 Non-Parametric Models
Survival models can be categorized into three groups: non-parametric, semi-parametric, and fully parametric (Wang et al., 2019). The group of non-parametric models comprises statistical estimators that provide information about data without any prior assumption on the event distribution. Non-parametric models rely on some notion of similarity between groups of individuals to improve the prediciton complexity. The most common non-parametric model is the Kaplan-Meier (KM) estimator (Goel et al., 2010), which is often used to plot the general survival behavior of a population. In fact, the KM estimator is not conditioned on the subjects’ features as it is tailored to provide aggregate information about the overall event distribution within the data. Another popular non-parametric model is the Random Survival Forest (Ishwaran et al., 2008), which builds a set of decision trees with the CART (Breiman et al., 1984) method by maximizing the event distribution difference between nodes according to repeated log-rank tests (Bland & Altman, 2004). Each leaf contains a non-parametric estimation of the subjects corresponding assigned to that specific terminal node. The final prediction is obtained by averaging the predictions of the trees in the forest.
@@ 2.2 Semi-Parametric Models
[…]
@@
Several extensions of the Cox model have been proposed, all relying on the proportional hazard assumption and partial log-likelihood optimization. Among those, CoxBoost (Ridgeway, 1999) and XGBoost (Chen & Guestrin, 2016) optimize the loss using gradient-boosted decision trees to estimate $r(\mathbf{x})$. On the other hand, DeepSurv (Katzman et al., 2018) replaces the linear dependency between parameters and features with a single-output neural networks.
@@ 2.3 Fully Parametric Models
Other popular models do not rely on predefined statistical distributions to construct survival estimations, but leverage neural networks to estimate the event probability directly at a fixed set of time intervals (Kvamme et al., 2019; Kvamme & Borgan, 2021). These neural networks have a single output per time bin, representing the event probability for that interval. One such model is DeepHit (Lee et al., 2018), a discrete-time survival model consisting of a shared feature extractor followed by event-specific sub-networks estimating the probabilities for each event. While these discrete-time models have shown promising practical results, they struggle with fine-grained or long-term prediction horizons due to their fixed-time nature. To address this limitation, some studies have proposed interpolation techniques between time points (Kvamme & Borgan, 2021; Archetti et al., 2024).
@@ 3.1 Hazard Function Definition
In FPBoost, the hazard function is composed of $J$ heads, each corresponding to a parametric distribution. This work includes heads following either the Weibull or LogLogistic distributions but, in principle, any hazard distribution differentiable with respect to its parameters can be included as a head. Table 1 collects the formulation of the considered distributions. This choice comes primarily from the fact that Weibull distributions are suited to model constant and increasing risks, while LogLogistic ones provide decreasing and arc-shaped risk profiles. Typically, the former is associated to survival behaviors such as aging and wear, while the latter relates to infant mortality. When combined, the resulting risk can assume a bathtub shape, covering a wide range of real-world behaviors (Klein et al., 2003; Nagpal et al., 2021). This approach provides an advantage from both a learning and interpretation perspective, as each sample is associated with a set of well-known distribution functions, where each parameter has a clear interpretation. In particular, for both the considered distributions, the $\eta$ parameter controls the scale – or spread – while the $k$ parameter changes the curve shape by modifying the slope or adding a bump.
@@
To guarantee the validity of the Weibull and LogLogistic hazard formulations, the scale and shape parameters $\eta_{j},k_{j}$ must be nonnegative. To enforce this constraint, we apply a ReLU activation function (Nair & Hinton, 2010) to these parameters, ensuring that any negative estimates are set to zero.
@@
[…]
@@
The proof, detailed in Appendix A.1, is a direct consequence of the Weierstrass Approximation Theorem, after showing that a single Weibull head is equivalent to a monomial of arbitrary degree. This underpins the capability of FPBoost to approximate any target hazard function with sufficient heads. To the best of our knowledge, this represents the first universal approximation theorem for machine learning models in the context of survival analysis. It is important to note that while only Weibull hazards are required for this theoretical result, incorporating LogLogistic hazards often enhances model performance and makes optimization easier in practical applications.
@@ 3.2 Loss Function and Training Procedure
In order to prevent overfitting, we add to $L_{\text{lik}}$ an ElasticNet regularization term (Zou & Hastie, 2005) as
@@ 4.1 Datasets
To ensure fair evaluation and consistency with similar studies, we selected datasets from well-known benchmarks in survival analysis, covering different conditions like breast cancer, lung cancer, AIDS, and cardiovascular diseases. Specifically, the AIDS (Hosmer et al., 2008), Breast Cancer (Desmedt et al., 2007), FLCHAIN (Dispenzieri et al., 2012), GBSG2 (Schumacher et al., 1994) and Veterans (Kalbfleisch & Prentice, 2011) datasets are provided by the scikit-survival (Pölsterl, 2020) Python library. The METABRIC (Katzman et al., 2018) and WHAS (Hosmer et al., 2008) datasets, instead, are available in the DeepSurv repository (Katzman et al., 2018) with a predefined test set. Lastly, SUPPORT2 (Knaus et al., 1995) is provided by SurvSet (Drysdale, 2022). Details on data collection and content overview are provided in Appendix A.2. Table 2 collects the summary statistics of these datasets.
@@ 4.2 Metrics
where $\mathbf{1}(\cdot)$ is an indicator function and $w_{i}(t)$ is a weighting factor that adjusts the censoring bias. This adjustment is the Inverse Probability of Censoring Weighting (IPCW) (Robins & Rotnitzky, 1992; Uno et al., 2011), which assigns weights based on the inverse probability of censoring at a given time $t$. The overall calibration of a survival model over time is summarized by integrating the Brier score across the entire study period, yielding the Integrated Brier Score (IBS).
@@ 4.3 Experimental Procedure
In order to validate FPBoost, we compared its performance against several baseline models. For these models, we employ default parameters provided by implementations in Python libraries. Specifically, we utilize scikit-survival for RSF, Cox, and CoxBoost, pycox (Kvamme et al., 2019) for DeepSurv and DeepHit, xgboost (Chen et al., 2022) for XGBoost, and auton-survival (Nagpal et al., 2022) for DSM. Following (Katzman et al., 2018), neural network architectures for DeepSurv, DeepHit, and DSM comprise three layers with neuron counts of 3, 5, and 3 times the number of features, respectively.
@@
The source code for these experiments is available at [Anonymized Link]. The FPBoost class implementation is fully compatible with scikit-survival, facilitating its inclusion into existing codebases.
@@ 5.1 Discussion and Future Work
While the idea of leveraging a mixture of parametric functions (Nagpal et al., 2021) has been previously explored in the survival literature, as well as ensemble learning (Ishwaran et al., 2008; Archetti et al., 2023), FPBoost introduces several innovations.
@@ A.1 Proofs
[…]
@@ A.2 Dataset Details
Figure 2: Kaplan-Meier estimations (blue) on survival probability and censoring probability (orange) for the datasets included in the study.
@@
The Veteran Administration Lung Cancer Trial (Veterans) (Kalbfleisch & Prentice, 2011) dataset focuses on lung cancer patients treated with two different chemotherapy regimens. This dataset is frequently used in simple survival benchmarks due to its small sample size.
@@
Figure 2 shows the Kaplan-Meier estimations on survival probability and censoring probability for the datasets included in the study.
@@ A.3 Hyperparameter Tuning
  • $\alpha$: Scaling factor for the ElasticNet loss. Search values are in $[0,\dots,1]$.
  • $\gamma$: Ratio between $L1$ and $L2$ penalties in ElasticNet. Search values are in $[0,\dots,1]$.
@@
  • Initialization: How to initialize parameter estimators $F_{\cdot_{j}}^{(0)}$. If Random, parameters are initialized as $\eta\sim\mathcal{N}(0.5,1)$, $k\sim\mathcal{N}(0,2)$, $w\sim\mathcal{N}(0,1)$. If KM, instead, parameter initialization is based on the Kaplan-Meier estimator. Specifically, a Weibull and a LogLogistic distribution are fitted to the KM estimator, obtaining $\bar{\eta},\bar{k}$ for both distributions. Then, head parameters are initialized as $\eta\sim\mathcal{N}\left(\bar{\eta},\bar{\eta}/10\right)$ and $k\sim\mathcal{N}\left(\bar{k},\bar{k}/10\right)$. In the KM case, weights are initialized uniformly.
@@ A.4.1 Results on C-TD and AUC
This section provides additional results on Time-Dependent Concordance index (C-TD) and AUC (Pölsterl, 2020). Specifically, the C-TD computes a concordance index weighting the contribution of each sample with the inverse censoring probability (IPCW) (Robins & Rotnitzky, 1992; Uno et al., 2011) following the same procedure as the IBS metrics. This reduces the bias in the concordance measure introduced by skewed censoring occurrences. Table 6 collects the experiment results on the C-TD metrics.
@@ A.4.2 The Impact of Model Type
Non-Parametric $+1.5$ $+1.5$
Semi-Parametric $+1.6$ $+6.4$ $+1.9$ $+4.1$
Fully Parametric $+7.3$ $+7.3$
All $+1.6$ $+4.8$ $+5.5$ $+4.6$
@@
Non-Parametric $-0.3$ $-0.3$
Semi-Parametric $-1.5$ $-1.2$ $-2.3$ $-1.7$
Fully Parametric $-5.6$ $-5.6$
All $-1.5$ $-0.7$ $-4.5$ $-2.8$

### v3

@@ 1 Introduction
In order to model the survival function, traditional methods often rely on simplifying assumptions, such as the risk proportion between different subjects being constant over time (Cox, 1972). These assumptions allow the construction of survival functions from a small set of parameters that can be estimated with statistical methods. While suited for contexts with limited data, these simplifying assumptions constrain generalization in real-world scenarios (Katzman et al., 2018). Machine learning techniques have allowed to advance these models by incorporating decision trees and neural networks, significantly enhancing their ability to identify and learn non-linear interactions within high-dimensional features. However, most of these techniques still operate under certain constraints, such as time discretization in neural-based approaches (Kvamme and Borgan, 2021) or accelerated failure time in tree-based gradient boosting (Collett, 2023).
@@
In this context, we introduce Fully Parametric Gradient Boosting (FPBoost), a novel architecture designed to model hazard functions through the composition of multiple fully parametric hazard functions. Hazard functions are related to survival functions as they measure the instantaneous risk of a subject experiencing the event of interest. FPBoost combines the strengths of tree-based ensemble learning with gradient boosting (Friedman, 2001; Collett, 2023), offering a flexible model with robust generalization capabilities and minimal assumptions. Modeling hazard as a weighted sum of multiple, fully-parametric functions, referred to as heads, allows FPBoost to be trained by maximizing the full survival likelihood (Wang et al., 2019). This, in turn, removes the need for simplified assumptions such as partial likelihood (Cox, 1972; Katzman et al., 2018) or discrete losses (Kvamme and Borgan, 2021). Additionally, the continuous nature of the learned survival functions ensures a fine-grained estimation of the event distribution, without requiring interpolation techniques (Archetti et al., 2023b, 2024). We theoretically show how this framework makes FPBoost a universal approximator of hazard functions, allowing it to learn in principle any target hazard, provided enough heads. Lastly, gradient-boosted trees applied to tabular data—the most common data format in survival applications—has proven to be still competitive against neural network techniques (Grinsztajn et al., 2022), as highlighted by our empirical results.
@@
We evaluate FPBoost in the right-censored, single-event setting, which is the most common application of survival analysis. The performance of FPBoost is benchmarked against state-of-the-art survival models, including both tree-based and neural network-based models (Chen and others, 2024). Performance metrics include the concordance index (Uno et al., 2011) to measure the discrimination capabilities of the model and the integrated Brier score (Graf et al., 1999), tailored for calibration. Our experiments demonstrate that FPBoost outperforms alternative models in both discrimination and calibration in the majority of cases and matches their performance when it does not.
@@ 2.1 Non-Parametric Models
Survival models can be categorized into three groups: non-parametric, semi-parametric, and fully parametric (Wang et al., 2019). The group of non-parametric models comprises statistical estimators that provide information about data without any prior assumption on the event distribution. Non-parametric models rely on some notion of similarity between groups of individuals to improve the prediciton complexity. The most common non-parametric model is the Kaplan–Meier (KM) estimator (Kaplan and Meier, 1958), which is often used to plot the general survival behavior of a population. In fact, the KM estimator is not conditioned on the subjects’ features as it is tailored to provide aggregate information about the overall event distribution within the data. Another popular non-parametric model is the Random Survival Forest (Ishwaran et al., 2008), which builds a set of decision trees with the CART (Breiman et al., 1984) method by maximizing the event distribution difference between nodes according to repeated log-rank tests (Bland and Altman, 2004). Each leaf contains a non-parametric estimation of the subjects corresponding assigned to that specific terminal node. The final prediction is obtained by averaging the predictions of the trees in the forest.
Finally, BoXHED (Wang et al., 2020; Lee et al., 2021; Pakbin et al., 2025) is a nonparametric boosting method for hazard estimation specifically designed to handle time-dependent covariates. Like FPBoost, BoXHED employs gradient boosting to minimize the survival likelihood, but extends beyond static covariates to accommodate time-varying covariate processes. The key difference lies in how the gradient is computed: BoXHED derives its gradient through a smooth convex representation of the likelihood function to accommodate time-varying trajectories, whereas FPBoost leverages the explicit parametric forms of its hazard functions, which promotes interpretable risk components corresponding to known distributions.
@@ 2.2 Semi-Parametric Models
$S(t|\mathbf{x})=\exp(-H(t|\mathbf{x}))=\exp\left(-\int\limits_{0}^{t}h(u|\mathbf{x})du\right)$ (1)
@@
Several extensions of the Cox model have been proposed, all relying on the proportional hazard assumption and partial log-likelihood optimization. Among those, CoxBoost (Ridgeway, 1999) and XGBoost (Chen and Guestrin, 2016) optimize the loss using gradient-boosted decision trees to estimate $r(\mathbf{x})$. On the other hand, DeepSurv (Katzman et al., 2018) replaces the linear dependency between parameters and features with a single-output neural networks.
@@ 2.3 Fully Parametric Models
Other popular models do not rely on predefined statistical distributions to construct survival estimations, but leverage neural networks to estimate the event probability directly at a fixed set of time intervals (Kvamme et al., 2019; Kvamme and Borgan, 2021). These neural networks have a single output per time bin, representing the event probability for that interval. One such model is DeepHit (Lee et al., 2018), a discrete-time survival model consisting of a shared feature extractor followed by event-specific sub-networks estimating the probabilities for each event. While these discrete-time models have shown promising practical results, they struggle with fine-grained or long-term prediction horizons due to their fixed-time nature. To address this limitation, some studies have proposed interpolation techniques between time points (Kvamme and Borgan, 2021; Archetti et al., 2024).
@@ 3.1 Hazard Function Definition
In FPBoost, the hazard function is composed of $J$ heads, each corresponding to a parametric distribution. This work includes heads following either the Weibull or LogLogistic distributions but, in principle, any hazard distribution differentiable with respect to its parameters can be included as a head. Table 1 collects the formulation of the considered distributions. This choice comes primarily from the fact that Weibull distributions are suited to model constant and increasing risks, while LogLogistic ones provide decreasing and arc-shaped risk profiles. Typically, the former is associated to survival behaviors such as aging and wear, while the latter relates to infant mortality. When combined, the resulting risk can assume a bathtub shape, covering a wide range of real-world behaviors (Klein et al., 2003; Nagpal et al., 2021). This approach provides an advantage from both a learning and interpretation perspective, as each sample is associated with a set of well-known distribution functions, where each parameter has a clear interpretation. In particular, for both the considered distributions, the $\eta$ parameter controls the scale (or spread) while the $k$ parameter changes the curve shape by modifying the slope or adding a bump.
@@
To guarantee the validity of the Weibull and LogLogistic hazard formulations, the scale and shape parameters $\eta_{j},k_{j}$ must be nonnegative. To enforce this constraint, we apply a ReLU activation function (Nair and Hinton, 2010) to these parameters, ensuring that any negative estimates are set to zero.
@@
Let $\mathcal{H}$ denote the space of hazard functions, that is, continuous nonnegative real functions $h(t)$ for which $\int_{0}^{\infty}h(t)\,dt=\infty$. For any $h^{\star}\in\mathcal{H}$, any $\varepsilon>0$, and any interval $[0,T]$, there exists a finite collection of $J$ Weibull hazard functions $h^{W}_{j}(t)$, with parameters $\eta_{j},k_{j}$, and weights $w_{j}$, such that
@@
The proof, detailed in Appendix A.1, is a direct consequence of the Weierstrass Approximation Theorem, after showing that a single Weibull head is equivalent to a monomial of arbitrary degree. This underpins the capability of FPBoost to approximate any target hazard function with sufficient heads. It is important to note that while only Weibull hazards are required for this theoretical result, incorporating LogLogistic hazards often enhances model performance and makes optimization easier in practical applications.
@@ 3.2 Loss Function and Training Procedure
In order to prevent overfitting, we add to $L_{\text{lik}}$ an ElasticNet regularization term (Zou and Hastie, 2005) as
@@ 4.1 Datasets
To ensure fair evaluation and consistency with similar studies, we selected datasets from well-known benchmarks in survival analysis, covering different conditions like breast cancer, lung cancer, AIDS, and cardiovascular diseases. Specifically, the AIDS (Hosmer et al., 2008), Breast Cancer (Desmedt et al., 2007), FLCHAIN (Dispenzieri et al., 2012), GBSG2 (Schumacher et al., 1994) and Veterans (Kalbfleisch and Prentice, 2011) datasets are provided by the scikit-survival (Pölsterl, 2020) Python library. The METABRIC (Katzman et al., 2018) and WHAS (Hosmer et al., 2008) datasets, instead, are available in the DeepSurv repository (Katzman et al., 2018) with a predefined test set. Lastly, SUPPORT2 (Knaus et al., 1995) is provided by SurvSet (Drysdale, 2022). Details on data collection and content overview are provided in Appendix A.2. Table 2 collects the summary statistics of these datasets.
@@ 4.2 Metrics
where $\mathbf{1}(\cdot)$ is an indicator function and $w_{i}(t)$ is a weighting factor that adjusts the censoring bias. This adjustment is the Inverse Probability of Censoring Weighting (IPCW) (Robins and Rotnitzky, 1992; Uno et al., 2011), which assigns weights based on the inverse probability of censoring at a given time $t$. The overall calibration of a survival model over time is summarized by integrating the Brier score across the entire study period, yielding the Integrated Brier Score (IBS).
@@ 4.3 Experimental Procedure
In order to validate FPBoost, we compared its performance against several baseline models. For these models, we employ default parameters provided by implementations in Python libraries. Specifically, we utilize scikit-survival for RSF, Cox, and CoxBoost, pycox (Kvamme et al., 2019) for DeepSurv and DeepHit, xgboost (Chen and others, 2022) for XGBoost, and auton-survival (Nagpal et al., 2022) for DSM. Following (Katzman et al., 2018), neural network architectures for DeepSurv, DeepHit, and DSM comprise three layers with neuron counts of 3, 5, and 3 times the number of features, respectively.
@@
The source code for these experiments is available at https://github.com/archettialberto/fpboost. The FPBoost class implementation is fully compatible with scikit-survival, facilitating its inclusion into existing codebases.
@@ 5.1 Discussion and Future Work
While the idea of leveraging a mixture of parametric functions (Nagpal et al., 2021) has been previously explored in the survival literature, as well as ensemble learning (Ishwaran et al., 2008; Archetti et al., 2023a), FPBoost introduces several innovations.
@@ A.1 Proofs
Let $\mathcal{H}$ denote the space of hazard functions, that is, continuous nonnegative real functions $h(t)$ for which $\int_{0}^{\infty}h(t)\,dt=\infty$. For any $h^{\star}\in\mathcal{H}$, any $\varepsilon>0$, and any interval $[0,T]$, there exists a finite collection of $J$ Weibull hazard functions $h^{W}_{j}(t)$, with parameters $\eta_{j},k_{j}$, and weights $w_{j}$, such that
@@ A.2 Dataset Details
Figure 2: Kaplan–Meier estimations (blue) on survival probability and censoring probability (orange) for the datasets included in the study.
@@
The Veteran Administration Lung Cancer Trial (Veterans) (Kalbfleisch and Prentice, 2011) dataset focuses on lung cancer patients treated with two different chemotherapy regimens. This dataset is frequently used in simple survival benchmarks due to its small sample size.
@@
Figure 2 shows the Kaplan–Meier estimations on survival probability and censoring probability for the datasets included in the study.
@@ A.3 Hyperparameter Tuning
  • $\alpha$: Scaling factor for the ElasticNet loss. Search values are in $0,\dots,1]$.
  • $\gamma$: Ratio between $L1$ and $L2$ penalties in ElasticNet. Search values are in $0,\dots,1]$.
@@
  • Initialization: How to initialize parameter estimators $F_{\cdot_{j}}^{(0)}$. If Random, parameters are initialized as $\eta\sim\mathcal{N}(0.5,1)$, $k\sim\mathcal{N}(0,2)$, $w\sim\mathcal{N}(0,1)$. If KM, instead, parameter initialization is based on the Kaplan–Meier estimator. Specifically, a Weibull and a LogLogistic distribution are fitted to the KM estimator, obtaining $\bar{\eta},\bar{k}$ for both distributions. Then, head parameters are initialized as $\eta\sim\mathcal{N}\left(\bar{\eta},\bar{\eta}/10\right)$ and $k\sim\mathcal{N}\left(\bar{k},\bar{k}/10\right)$. In the KM case, weights are initialized uniformly.
@@ A.4.1 Results on C-TD and AUC
This section provides additional results on Time-Dependent Concordance index (C-TD) and AUC (Pölsterl, 2020). Specifically, the C-TD computes a concordance index weighting the contribution of each sample with the inverse censoring probability (IPCW) (Robins and Rotnitzky, 1992; Uno et al., 2011) following the same procedure as the IBS metrics. This reduces the bias in the concordance measure introduced by skewed censoring occurrences. Table 6 collects the experiment results on the C-TD metrics.
@@ A.4.2 The Impact of Model Type
Non-Parametric +1.5+1.5 +1.5+1.5
Semi-Parametric +1.6+1.6 +6.4+6.4 +1.9+1.9 +4.1+4.1
Fully Parametric +7.3+7.3 +7.3+7.3
All +1.6+1.6 +4.8+4.8 +5.5+5.5 +4.6+4.6
@@
Non-Parametric 0.3-0.3 0.3-0.3
Semi-Parametric 1.5-1.5 1.2-1.2 2.3-2.3 1.7-1.7
Fully Parametric 5.6-5.6 5.6-5.6
All 1.5-1.5 0.7-0.7 4.5-4.5 2.8-2.8